### v2

# Macaw: A Machine Code Toolbox for the Busy Binary Analyst

Abstract.
When attempting to understand the behavior of an executable, a binary analyst can make use of many different techniques. These include program slicing, dynamic instrumentation, binary-level rewriting, symbolic execution, and formal verification, all of which can uncover insights into how a piece of machine code behaves. As a result, there is no one-size-fits-all binary analysis tool, so a binary analysis researcher will often combine several different tools. Sometimes, a researcher will even need to design new tools to study problems that existing frameworks are not well equipped to handle. Designing such tools from complete scratch is rarely time- or cost-effective, however, given the scale and complexity of modern instruction set architectures.
We present Macaw, a modular framework that makes it possible to rapidly build reliable binary analysis tools across a range of use cases. Over a decade of development, we have used Macaw to support an industrial research team in building tools for machine code–related tasks. As such, the name “Macaw” refers not just to the framework itself, but also a suite of tools that are built on top of the framework. We describe Macaw in depth and describe the different static and dynamic analyses that it performs, many of which are powered by an SMT-based symbolic execution engine. We put a particular focus on interoperability between machine code and higher-level languages, including binary lifting from x86 to LLVM, as well verifying the correctness of mixed C and assembly code.
1. Introduction
Binary analysis refers to not one class of tool, but many. Starting with an unknown binary, there are many tasks that an engineer may want to accomplish. These may include discovering the structure of the binary, transforming the binary to achieve some goal, analyzing the binary to identify particular properties or vulnerabilities, and/or formally verifying that the binary matches some specification. These broad tasks can be combined and specialized into many domains, objectives, tool types, and instruction sets. The result is that there is as yet no optimal binary analysis tool, but rather a large space of potentially useful designs.
We present Macaw, a binary analysis toolkit which we have built to help us explore this design space of tools. Macaw exists in a similar niche as frameworks such as Angr (Shoshitaishvili et al., 2016) and, to a lesser extent, reverse-engineering suites such as Ghidra (Rohleder, 2019) and Binary Ninja (Vector35, 2023a). What sets Macaw apart are the intended use cases. Macaw is not a standalone tool, but it is instead a Haskell library that can be used to rapidly construct and evaluate binary analysis tools. The main Macaw library is also accompanied by a suite of libraries for disassembly, representing architecture semantics, and symbolically executing machine code, which we collectively refer to as the Macaw ecosystem. Macaw is designed to help us build quickly, maximize reuse of existing components, and avoid costly errors during prototyping. This imposes a particular set of design constraints on the Macaw ecosystem, which we explain in this paper.
Macaw was built to support an industry research team developing novel binary analysis tools. This team executed on multiple research projects over a decade, with a significant amount of personnel change over this time. In this environment, Macaw has served as a key library that helped us develop successful tool prototypes. In this paper, we discuss the most interesting and mature tool designs, which are:
Reopt, a binary lifter which converts x86-64 binaries to LLVM code and performs reoptimizations on the lifted code. This was the original Macaw use-case.
Renovate, a static binary rewriter that allows users to add or remove code without executing the binary.
MCTrace, a binary instrumentation tool to insert probes to dynamically collect telemetry with low-overhead tracing, including in environments without an operating system.
Ambient-Verifier, a static verifier that can prove the presence of weird machines (Dullien, 2020) in binaries (or lack thereof).
PATE, a relational verifier that proves that two binaries have the same observable behaviors.
Cerridwen, a tool for ranking how similar a binary is to a known corpus of binaries.
Surveyor, an interactive debugger that, among other features, can step through machine code and inspect symbolic values.
To illustrate in more depth how Macaw works, we discuss two of these tools in detail, Reopt and SAW. We focus on these two because they support very different tasks and, as such, require combining Macaw’s components in very different and illustrative ways.
A key component of Macaw’s design is its deep integration with Crucible (Christiansen et al., 2019), an SMT-backed symbolic execution engine. Although Crucible was originally developed for simulating higher-level imperative programming languages such as C, the same technology proves valuable for simulating the behavior of machine code as well. Many tools built on top of Macaw—including SAW, Renovate, Ambient-Verifier, PATE, and Surveyor—leverage Crucible to peform different analyses, including static program analysis and formal verification. Because Crucible also supports higher-level languages such as C, this makes it easier to simultaneously analyze codebases that mix these languages with machine code. This is something that SAW takes advantage of, as it leverages Macaw and Crucible to verify mixed C and machine code.
Macaw provides a specific point in the design space of binary analysis frameworks which has proved useful for us in developing a wide range of capabilities. In Section 2, we discuss the core design of the Macaw IR and supporting ecosystem, and we discuss how this design aligns with goals. In Section 3, we explore the tools that we have built on top of Macaw in more detail. We use this to explain how Macaw helps us to effectively build tools in a variety of domains. In Section 4, we compare Macaw to existing frameworks such as Angr, and we also compare Macaw-based tools to similar efforts. Many of the differences come down to tradeoffs; we describe both the upsides and downsides of the choices that led to Macaw’s design. Section 5 concludes and discusses our future design ambitions for Macaw.
2. Design of the Macaw Ecosystem
The Macaw ecosystem consists of a set of Haskell libraries, whose relationships are depicted in Figure 1. At the heart of the ecosystem is Macaw-Base, which only contains functionality that is independent of any particular architecture. This includes the core Macaw intermediate representation (IR), code discovery, optimization passes, and an ELF loader for ingesting binaries.
Any architecture-specific functionality is encapsulated in a Macaw backend library. Macaw was originally designed with x86-64 support in mind, which gave rise to the first backend, Macaw-x86. Later, Macaw-AArch32, Macaw-PPC, and Macaw-RISCVwere added. Each of these backends are built on top of other libraries that integrate Macaw with disassemblers for specific instruction set architectures (ISAs), and they also encode the semantics of each architecture’s instructions into a form that can be reasoned about. For instance, Macaw-RISCV is built on top of the GRIFT library, which provides a formal specification of the RISC-V ISA in Haskell (Selfridge, 2019).
2.1. Integrating ISAs into Macaw
While Macaw-Base provides the basic substrate for performing binary analysis, it is not very useful without a mechanism for representing ISA-specific details within Macaw. To this end, there are a suite of libraries supporting each architecture backend. Each Macaw-supported architecture comes with its own disassembler for parsing machine code into a form that Macaw can ingest, and it also comes with a library for representing the semantics of each instruction in a precise way. Having codified semantics enables both formal verification, which Macaw richly supports (see Section 3.2), and machine-code synthesis (Srinivasan and Reps, 2015), for which Macaw has experimental support.
Incorporating a new ISA into Macaw can be a substantial amount of work. Throughout the years, we have experimented with a number of different approaches for implementing Macaw disassemblers and semantics, and we have made pragmatic choices over time. For some architecture backends, such as Macaw-RISCV, we implement disassemblers and semantics by hand. For other backends, we generate implementations automatically from ISA specifications. For instance, Macaw-AArch32’s disassembler is derived from the official ARM XML specification (Reid, 2016), which we use to compute parse tables for each instruction. In addition, the XML spec encodes the semantics of each instruction in ARM’s Architecture Specification Language (ASL). We generate Macaw-AArch32’s semantics by symbolically executing ASL expressions using Crucible (see Section 2.6).
Each approach has its tradeoffs. Manually implementing disassemblers and semantics makes it easier to understand the code in isolation, but it requires more careful auditing for mistakes. Machine-generated code is generally quicker to develop and less error-prone, but at the expense of obfuscating the code. As a general principle, we prefer hand-written implementations for smaller ISAs (e.g., RISC-V) and machine-generated implementations for larger ISAs (e.g., AArch32).
2.2. Typed Address Representations
Macaw represents a machine address with the MemWord w type, which consists of a machine word of size w bits. Note that w is a type-level number, which means that this convention is checked at compile-time, not runtime. This makes critical use of GHC’s DataKinds extension, which allows certain forms of values to be promoted to the type level (Yorgey et al., 2012).
The w type parameter is one example of a key design choice made throughout Macaw’s code: where possible, use GHC language extensions to encode invariants about machine code at the type level. This tradition was inspired by the Crucible symbolic execution library (Christiansen et al., 2019), with which Macaw integrates (see Section 2.6). For example, (Christiansen et al., 2019, Section 3) defines a CrucibleType data type that classifies the shape of Crucible expressions. Following suit, Macaw defines its own MacawType to classify the shape of machine code values—we give some selected definitions in Appendix A, Figure A.1.
An example of this strong typing discipline being put to use is encoding the bit widths of AArch32 instructions. AArch32 processors have two instruction sets: ARM mode, where each instruction is 32 bits, and Thumb mode, where each instruction is 16 bits. Macaw encodes these bit widths at the type level, and as a result, attempting to mix the two instructions sets without an explicit mode switch will result in a type error. This has caught many potential errors when developing Macaw-based tools, as ARM and Thumb instructions can often occur within the same binary.
2.3. From Addresses to Architectures
Just as Macaw’s address types are indexed by the number of bits w, many Macaw abstract syntax types are parameterized by an arch type parameter, which represents the processor architecture being used. Doing so ensures that operations from separate architectures are not mixed up. For example, Macaw-x86 instantiates arch with an X86_64 type, Macaw-AArch32 instantiates arch with an AArch32 type, and so on. We refer to types like X86_64 and AArch32 as architecture extension points. One example of an arch-indexed type is ArchReg, which is displayed in more detail in Appendix A, Figure A.2. ArchReg uses GHC’s TypeFamilies extension (Schrijvers et al., 2008) to map an architecture extension point to a register data type.
Sometimes, arch-indexed data types must interact with w-indexed data types. A Macaw architecure uniquely determines the number of bits in a machine address, so Macaw defines a RegAddrWidth type family to bridge the gap between the former and the latter. This is such a commonly used type family that Macaw also provides a variety of type synonyms defined on top of RegAddrWidth for convenience, which are depicted in Appendix A, Figure A.2.
2.4. The Macaw Intermediate Language
Macaw uses a three address code–based intermediate language (Aho et al., 1986) that is centered around basic blocks (Block). Each basic block consists of a list of statements (Stmt) followed by a special terminator statement (TermStmt). These data types are given in Appendix A, Figure A.3. For example, the Block type is defined as follows:
Each data type is parameterized by two type parameters: arch, which encodes the architecture, and ids, which denotes the set of identifiers used in assignment instructions. The ids type parameter is unique to each decoded function, which ensures that instructions from different functions cannot be mixed without an explicit renaming step. The ids type parameter functions similarly to the s type parameter in Haskell’s ST s type (Launchbury and Peyton Jones, 1995).
The control flow in Macaw is oriented around basic blocks, and each basic block has an address corresponding to the address of the first instruction. A disassembled machine instruction can correspond to multiple Stmts, depending on the complexity of the instruction. Note that the Stmts in two blocks are allowed to overlap. For example, this is necessary to support AArch32 binaries, where instructions can be in both ARM or Thumb mode depending on how they are reached.
The core Macaw IR is compact and covers common operations present in multiple ISAs, which are used to represent the behavior of many instructions. These operations primarily involve integer arithmetic (e.g., add and sub in x86-64), and Macaw encodes these operations using bitvectors in a way that can be easily translated to SMT. More complicated operations (i.e., ones that do not map directly to SMT) are put into syntactically separate parts of the Stmt language. For example, memory accesses are their own category, as they must be decomposed into SMT-friendly operations via symbolic execution (see Section 2.6). There is also a separate category for architecture-specific extensions (e.g., data cache hints in PowerPC), for which each ISA has a small number of instructions.
This separation between common operations and more complicated ones reflects a key Macaw design decision: the Macaw IR does not attempt to be generic enough to encode every possible instruction directly. One benefit for this decision is that it keeps the Macaw IR relatively readable. Although Macaw could expand every instruction directly to core operations, this could easily cause the number of Stmts to balloon in size. For example, x86-64’s AES-NI instructions would likely need hundreds of bitvector operations to encode directly. Another benefit is that some instructions perform effects that not all Macaw clients need to model, so putting them into architecture-specific extensions means that clients only “pay” for extensions that they reason about. This gives Macaw users a high degree of control when mixing different components.
2.5. Code Discovery
One of the most important passes in Macaw-Base is code discovery, which aims to identify all of the functions in a binary that are reachable from a set of entry point addresses. The core algorithm that powers code discovery is illustrated in Figure 2. The algorithm discovers one function at a time by taking candidate function addresses from the exploration frontier, which is seeded with an initial set of entry points. Usually, these entry points are taken from the entry point address of an ELF binary, the dynamic entry points of a shared library, the addresses of static symbols (when available), as well as function addresses derivable from metadata, such as DWARF unwinding tables. Macaw also supports specifying custom entry points, which can be useful for binaries that lack symbol information (e.g., stripped binaries).
The main part of the algorithm is the block decoding loop, which discovers functions by decoding their individual blocks. First, Macaw decodes an instruction using a low-level disassembly function for the specific architecure. Next, Macaw lifts the instruction into a semantic representation as a sequence of Stmts. If the instruction changes control flow (e.g., x86-64’s call), Macaw will terminate the block with a TermStmt; otherwise, it will restart the loop and continue decoding the block.
Next, Macaw performs rewriting, which simplifies terms in the block to improve readability and simplify later analyses. Afterwards, Macaw analyzes the block to determine what kind of control-flow transfer terminates the block, be it a jump, function call, or otherwise. To this end, Macaw tracks abstract domains for each machine register, using techniques inspired by value-set analysis (Balakrishnan and Reps, 2004) to update the domains as new blocks are discovered. At the end of analysis, Macaw consults the abstract return address, as well as the addresses and bounds of jump tables, to classify how the block terminates. Targets of function calls are added to the exploration frontier as they are classified, which is used in subsequent iterations of the algorithm.
2.6. Symbolic Execution via Crucible
A key component of the Macaw ecosystem is Macaw-Symbolic, which defines functionality for simulating Macaw IR programs with Crucible, a symbolic execution library targeting imperative code (Christiansen et al., 2019). Using Macaw-Symbolic, one can simulate a binary with symbolic data, such as symbolic register states or stack contents. During simulation, Crucible will generate verification conditions that are discharged to an SMT solver.
Besides machine code, Crucible also includes backends targeting other imperative programming languages, including C, Java, and Rust. Crucible’s LLVM backend (which enables reasoning about C code) has been battle-tested on a large number of verification problems from the SV-COMP verification competition (Scott et al., 2021). Moreover, Crucible can reason about code that mixes languages, such as C programs that use inline assembly (see Section 3.2).
2.6.1. Symbolic Code Discovery
The Macaw ecosystem also includes an additional, optional method of code discovery provided by the Macaw-Refinement library. This library utilizes symbolic execution to discover code targets, which can supplement the simpler, pattern-based heuristics that Macaw-Base uses.
As an example, consider a switch statement whose expression is an integer i whose range of possible values was previously constrained by an if statement. When the various case targets of the switch statement are small and regularly sized, the compiler may encode the switch statement as a jump to i * s, where s is the size of the largest case block. Macaw-Base’s pattern-based heuristics will not be aware the constraints from the if statement, and thus they cannot limit the scope of possible jump targets to the specific set of valid addresses. Using Macaw-Refinement, on the other hand, allows an SMT solver to indicate the constrained set of possible jump targets to add those to the discovery frontier.
3. Case Studies
3.1. Reopt
Reopt (Hendrix et al., 2019) is an optimization tool that lifts functions in a compiled binary into LLVM bitcode, optimizes the intermediate bitcode, and then recompiles it into a new executable. Reopt currently supports x86-64 binaries compiled in the ELF format for Linux. The Macaw library began its life as an internal component of the Reopt tool, and much of Macaw’s design reflects considerations that were important at the time of Reopt’s development.
The primary challenge in designing a tool like Reopt is going from Macaw’s IR, which only has very minimal type information associated with each basic block, to the LLVM IR, which has a comparatively richer type system. At the Macaw level, most values in general-purpose registers look like integers, but when Reopt outputs LLVM bitcode, it is helpful to distinguish between integers, structs, pointers, and function pointers.
To this end, Reopt includes a type-based function recovery pass, which is depicted in Figure 3. Starting from a list of initial entry point functions (which are either inferred from the binary or user-specified), the tool calls Macaw to discover the closure of basic blocks reachable from the currently known blocks. Once the basic blocks have been recovered (c.f. Section 2.5), Reopt performs a demand analysis, allowing it to compute the x86-64 arguments and return registers on a per-block basis, aggregating them into register “demand sets” for entire functions, which are then synthesized into appropriate argument and return values for the whole function.
At this point, we only know the size of registers, but not whether they contain integers, floating-point values, or, for those registers whose size matches that of virtual addresses for the architecture, data or code pointers. However, there is information in the instruction flow that can help distinguish between all these cases, sometimes uniquely identifying what high-level type a register should have. In order to produce better LLVM code, Reopt will gather constraints for all intermediate registers, function arguments, and return values. For instance, a register flowing into the memory operand of a memory operation ought to be some pointer, while a value coming out of a division operation must be numeric. Reopt will solve these constraints to determine a candidate type for every function. Reopt also has some experimental support for row types, allowing it to recover structural types, including recursive ones such as linked lists.
During this process, Reopt may uncover calls to other functions, in which case Macaw’s code discovery will need to be invoked once more. For instance, if a function’s argument is determined to hold function pointers, all concrete values flowing into this argument position can be safely added to the candidate entry points pool. Thanks to Macaw’s incremental design, Reopt is able to resume block discovery with those new entry points without incurring duplicate work. Once all of the constraints are solved, the final output is an LLVM program consisting of all the function definitions that have been recovered, which now have much richer types.
3.1.1. Reopt Type Constraints
Reopt’s approach to reverse engineering types for assembly functions is loosely based on the TIE system described in Lee et al. (2011). The type system presented in that work is slightly more sophisticated than what Reopt needs. For instance, Reopt generates LLVM, which does not distguish between the types of signed and unsigned integers, so Reopt does not bother to track this information. Likewise, where TIE relies on value-set analysis to get more precise offset information in structured types, Reopt relies on Macaw’s symbolic offsets, which may either be concrete offsets or opaque to its analysis.
Ultimately, Reopt is concerned with distinguishing pointers from non-pointer values. Without the type inference pass, Reopt would lift all pointer arithmetic to integer arithmetic and would have to introduce an LLVM inttoptr cast every time such a value would be in address position in a memory load or store instruction. This is important because treating pointers as integers up to the last second inhibits LLVM optimizations and obfuscates pointer manipulation for both human and tool consumers of the lifted IR. Instead, pointer arithmetic can be lifted to pointer-manipulating instructions, describing them using types with richer structure, benefiting the readability of the lifted code, and enabling more complex code analyses and transformations. Knowing these types, Reopt can produce the more appropriate getelementptr LLVM instruction, operating over typed pointers. We give an instance of such an output change in Appendix B, Figure B.1.
3.1.2. Reopt Limitations
Reopt’s function recovery, as sophisticated as it is, may still not be enough to infer a precise type for a function. One reason is that Macaw’s code discovery is not perfect, and when it cannot process all of the instructions in a basic block, it will conservatively bail out and possibly omit some instructions that are crucial for inferring the types of the block’s inputs and outputs. To help alleviate this problem, Reopt includes an option to supply header files, as well as debug information files, which contain function type signatures that Reopt should assume when it encounters known functions. This is particularly useful for external function calls (e.g., from libc), where Reopt either does not have the entire function definition available, or the full definition is sufficiently low-level that analyzing it proves costly.
Recent work has also been aimed at addressing limitations when working with stripped binaries with no debug sections. Such binaries have historically proven challenging for Reopt, as the only entry point initially known for such binaries is the one listed as entry point of the executable. Typically, this will be the address of a symbol that would have been named _start before stripping, and whose code does not contain an explicit call to the address of the main symbol.
Instead, the start routine will call the C run-time scaffolding symbol, __libc_start_main, passing the address of the main symbol as an argument to it. As such, the binary does not contain any explicit call to main, and Macaw will subsequently miss the main function and stop exploring right away. This is where the type-driven analysis of Reopt can help. In this case, Reopt knows that __libc_start_main expects a function pointer thanks to the libc debug information. Upon noticing the address of main flowing into a call to __libc_start_main, Reopt will instruct Macaw to resume block discovery with this new address as a function entry point.
There are three reasons why we designed Reopt to use a type-based approach. First, Reopt favors this approach over heuristical ones, as this not only applies to discovering the main symbol in a stripped binary, but also any function address passed to a higher-order function, be it in an external library or locally within the binary being lifted. Second, Reopt prefers a discovery-guided approach, where we only inspect bytes of the code section when we witness an explicit flow of control to it, over a greedy inspection of the entire code section, as it still allows us to avoid false positives such as dead code, nop paddings, or misaligned decodings. Finally, Reopt favors a type-aware approach to a greedy inspection of all pointer-sized values that could look like they could point to code, as some floating-point values may also look like pointers to the code segment, and this would also yield false positives. In all the above cases, not only does Reopt find known-to-be-reachable addresses, it also discovers them in a context where it knows the type of their arguments and return value, which allows Reopt to properly lift these basic blocks into well-typed LLVM functions.
3.2. SAW
The Software Analysis Workbench (SAW) (Dockins et al., 2016) is a verification tool that can reason about the behavior of imperative programs by proving them equivalent to functional specifications. SAW supports a variety of different source languages, including code written in C, Java, Rust, and most relevantly for this work, x86-64. SAW has been successfully used to verify a variety of industrial-strength cryptographic libraries written in a mixture of C and x86-64, including the s2n Transport Layer Security (TLS) stack (Chudnov et al., 2018), the AWS LibCrypto library (Boston et al., 2021), and the BLST library (Dodds, 2020).
SAW ingests x86-64 machine code by leveraging Macaw to lift assembly instructions into a Crucible control-flow graph. Separately from the machine code itself, SAW also takes as input a list of specifications that describe the intended behavior of each assembly function. These specifications consist of Hoare-style pre- and post-conditions that describe the shape of memory and the behavior of the function itself. SAW checks each function against its specification by symbolically executing the function’s CFG and ensuring that the conditions described in the function are respected along the way, discharging any non-trivial proof goals to an SMT solver.
Aside from checking functional correctness, SAW also checks that a function is memory safe. In the context of a language like C, this is a well understood problem, as the C standard dictates what it means for a C program to interact with memory in a valid way. The notion of memory safety becomes blurrier with assembly code, however, as there are almost no universally accepted rules on what constitutes safe reads and writes. The only guidelines that one can typically rely on are the conventions used within a particular program, as well as the Application Binary Interface (ABI) for calling external functions.
SAW specifically targets programs with mixed C and x86-64 code where the C code controls the program’s memory, and all x86-64 functions only ever access the C code’s memory. In this setting, SAW can leverage Crucible’s LLVM-based memory model for modeling both types of code. This means that SAW is able to prove all memory reads and writes in the x86-64 code are correct with respect to the host language’s memory model. As a small example of how SAW’s x86-64 support works, we will verify the behavior of a tiny increment assembly function that takes a single pointer argument and increments the value it points to by one.
We can write a SAW specification increment_spec that captures the intended behavior:
let increment_spec = do {
ptr <- llvm_alloc (llvm_int 64);
val <- llvm_fresh_var "val" (llvm_int 64);
llvm_points_to ptr (llvm_term val);
llvm_execute_func [ptr];
llvm_points_to ptr (llvm_term {{ val + 1 }});
This specification declares a series of statements, where each statement before the llvm_execute_func is a pre-condition and each statement after is a post-condition. The pre-conditions state that the function accepts a single pointer argument that points to an arbitrary 64-bit LLVM integer value. The post-condition states that after the function is invoked, the pointer argument will then point to the initial integer value plus one. Again, we are leveraging LLVM to describe the shape of memory, and all of the commands used here could apply to C functions just as much as x86-64 ones.
The increment function can be checked against increment_spec by running:
llvm_verify_x86 "./increment_bin" "increment" increment_spec z3;
Here, increment_bin is the ELF binary that defines increment, and z3 is the SMT solver to discharge proof goals to during symbolic execution. This is the only command that is machine code–specific, as this is where Macaw is used to lift the assembly code in increment to a CFG and pass it on to SAW’s symbolic execution engine. Although the llvm_verify_x86 command is specific to the x86-64 ISA, it could very easily be extended to handle other architectures as well 111For example, an llvm_verify_aarch32 command is planned: anonymized for review .
3.3. Other Case Studies
Ambient-Verifier
A verification tool built on top of Macaw that evaluates whether or not it is possible to trigger a weird machine (WM) (Dullien, 2020) in a binary, i.e., a code execution path that occurs outside of the intended specification of the program. One challenging aspect of triaging the presence of WMs is determining what environmental conditions—such as command-line arguments, environment variables, or file system state—are strictly required to trigger the WM, and which conditions are irrelevant to the WM. To this end, Ambient-Verifier leverages symbolic execution to make certain parts of the environment symbolic to explore the different execution paths, which can help users of the tool determine which parts of the environment are most relevant to the WM at hand.
Many WMs reside in a shared library that a binary depends on, such as libc. To this end, it is important to be able symbolically execute the code inside of low-level shared libraries, but this can be difficult due to how often these libraries contain low-level functions and system calls. While Crucible can simulate this type of code, the performance is often impractically slow.
To allow the tool to execute shared libraries in an efficient manner, Ambient-Verifier includes a mechanism for overriding the behavior of functions and system calls. For instance, Ambient-Verifier can intercept calls to the getppid system call and replace it with a version that returns a fresh, symbolic integer as the process ID. While this does not faithfully emulate the actual implementation of getppid, this can be an acceptable compromise to make to speed up symbolic execution, especially for WMs where the behavior of getppid is irrelevant.
PATE (Patches Assured up to Trace Equivalence)
A relational verification (Barthe et al., 2011, 2013) tool that proves that a property about a binary’s observable behavior holds before and after a patch is applied to the binary’s code. A typical use case for PATE is verifying that security-oriented patches fix undesirable behaviors in a binary without adversely affecting other parts of the binary. First, PATE splits the original and patched binaries into conflict-free acyclic regions (CFARs). A CFAR is a collection of basic blocks with control flow between them, but with no backedges. PATE uses the original binary as the behavioral specification, inferring frame conditions for each CFAR. PATE then symbolically executes each CFAR in the original and patched binary, and if the patched binary fails to satisfy the frame condition inferred from the original binary, then PATE produces a summary of the conditions that lead to exhibiting different behavior.
Cerridwen
A tool that quantifies similarities among x86-64 binaries. The algorithm for ranking similarity largely follows the description in (David et al., 2017). Cerridwen decomposes a binary into strands—that is, data-flow slices of basic blocks (David et al., 2016)—which are used as units of measurement for comparing how similar the binary is to a corpus of known binaries. The basic blocks resulting from Macaw’s code discovery are used to compute each strand, which consists of the operations that lead to the computation of a single value within a basic block.
Renovate
A static binary rewriting library for ELF binaries. Using Renovate, one can add, remove, or rewrite basic blocks in a statically linked binary without needing to execute it. The core of Renovate’s rewriting machinery is architecture-independent, and there exist x86-64, PowerPC, and AArch32 backends. One particular use case for Renovate is achieving binary diversity (Cohen, 1993)—that is, randomizing the layout of basic blocks to make it more resilient to attackers.
MCTrace
An x86-64 and PowerPC binary instrumentation tool that is heavily inspired by DTrace (Cantrill et al., 2004; Gregg and Mauro, 2011), but does not require any special kernel support or recompiling the program to be traced. MCTrace reads the same tracing specifications as DTrace (written in the D language) and leverage’s Renovate’s binary rewriting capabilities to insert probes into binaries. The traces that these probes generate can support a variety of different systems. For example, MCTrace-instrumented binaries have been successfully run on bare metal using an MPC5777C PowerPC microcontroller.
Surveyor
A library for interactively debugging LLVM, JVM, x86-64, and PowerPC programs, using an interface inspirted by gdb and emacs. Notably, Surveyor is a symbolic debugger: it supports setting breakpoints during Crucible’s symbolic execution, and it can inspect the values of arguments to a basic block, which may possibly be symbolic themselves. Surveyor also supports stepping through a binary’s execution and recording a trace of symbolic events, which can be replayed later.
4. Related Work
Reverse Engineering Frameworks
The closest tools to Macaw in terms of capabilities are industrial binary analysis frameworks such as Ghidra (Rohleder, 2019), Binary Ninja (Vector35, 2023a), Radare2 (Alvarez, 2023), and IDA Pro (Eagle, 2011). Like Macaw, these frameworks handle a multitude of binary analysis tasks such as reverse engineering, disassembly, decompilation, and static analysis. Each framework also includes a native IR that abstracts some low-level architectural details—for example, Ghidra’s P-Code (Naus et al., 2023) and others (Vector35, 2023b; Radare2 Book, 2023; Rolles, 2018).
The main difference between these tool and Macaw lies in their respective design goals. Ghidra and similar are intended for industrial-strength reverse engineering applications. Macaw is intended to be a research-oriented toolkit for rapidly building novel binary analysis tools. For the user, these binary analysis frameworks offer a polished experience which Macaw does not seek to provide. For example, Ghidra features a full, graphical IDE that allows users to visualize disassembled programs, label fields, rename variables, and more. We have experimented with different UI approaches via the Surveyor tool, but as exploratory tools, not fully-developed capabilities.
Macaw’s greatest strength is providing a platform to build new binary analysis tools from scratch. Frameworks such as Ghidra typically support plugins via some API, but this is usually meant for extending the existing functionality rather than building standalone applications. In this sense, Macaw occupies a different niche in the design space. Indeed, it is possible to build a custom analysis tool with Macaw and integrate its capabilities into Ghidra’s IDE.
Another difference is Macaw’s intrinsically typed IR, which encodes invariants about well-formedness at the type level. This guarantees that IR values are valid in the context of the analyses they want to perform. To our knowledge, none of the other IRs mentioned earlier employ a comparable approach to IR validity. There have been efforts to give formal semantics to P-Code, but this effort is external to the actual P-Code representation used in Ghidra (Naus et al., 2023).
Binary Lifting
Reopt is one tool in the large space of binary lifters and rewriters; see Schulte et al. (2022) for a comparison. Reopt lifts machine code to the LLVM IR, which puts it in the same category as lifters such as McSema (Dinaburg and Ruef, 2014), SecondWrite (Anand et al., 2013), RevGen (Chipounov and Candea, 2011), and rev.ng (Di Federico et al., 2017). It differs primarily in the style of LLVM code that it produces. Most LLVM-based lifters map processor registers directly to LLVM variables with an explicit stack, which results in atypical LLVM code. Reopt unifies the machine code stack with the LLVM stack in an effort to produce more “human-friendly” LLVM.
Symbolic Execution of Machine Code
Macaw leverages Crucible’s support for forward symbolic execution (King, 1976; Schwartz et al., 2010) to analyze the behavior of programs. Other binary analysis frameworks use symbolic execution as well, examples of which include Angr (Shoshitaishvili et al., 2016), BAP (Brumley et al., 2011), Triton (Saudel and Salwan, 2015), Mayhem (Cha et al., 2012), and KLEE-Native (Vegasena and Goodman, 2019). Most of these leverage dynamic symbolic execution (also known as concolic execution), while Macaw uses static symbolic execution by lifting a binary to a high-level IR and analyzing it using symbolic values.
This reflects a difference in the types of problems we have typically targeted with Macaw. Other tools typically use symbolic execution in bug-finding (for which dynamic symbolic execution is well-suited) while we have often used Macaw to support exhaustive exploration or formal verification (for which static symbolic execution is necessary).
Angr is static insofar as it lifts binaries to an IR and performs additional analyses, but it is also provides dynamic analyses such as a symbolic-assisted fuzzer (Stephens et al., 2016). This is similar to our Macaw-based tool Ambient-Verifier, but Angr uses a combination of static and dynamic symbolic execution depending on the use case. For example, Angr can directly interact with files and the operating system, whereas Ambient-Verifier can only model filesystem and OS interactions in an abstract way.
Another important distinction is that Angr often concretizes certain types of symbolic values during execution, depending on the concretization strategy that the user chooses. With the default concretization strategy, Angr will concretize symbolic addresses that are written to, as well as symbolic length values passed to the read() and write() system calls. This is done primarily for performance reasons, as preserving all symbolic values can often lead to VCs that are prohibitively expensive to solve for. By contrast, Ambient-Verifier (and Macaw more generally) propagates symbolic values throughout its analyses. This means that Ambient-Verifier will explore more paths by default, but perhaps at the expense of additional analysis time.
Formal Verification of Machine Code
Our Macaw-based verification tool SAW is similar to Vale (Bosamiya et al., 2020; Fromherz et al., 2019; Bond et al., 2017), a tool for writing verified x86-64 assembly code. Vale integrates with Low* (Protzenko et al., 2017) to allow verification of mixed C/x86-64 code. Vale and Low* are used in the implementation of EverCrypt (Protzenko et al., 2020), a library of verified C/x86-64 cryptography.
Unlike Vale and Low*, SAW is primarily used to verify pre-existing C/x86-64 code, whereas Vale and Low* are primarily meant for writing code and proofs simultaneously. Because SAW verifies x86-64 code as-is, SAW makes essential use of Macaw’s binary lifting as part of the overall verification process. Vale, on the other hand, generates x86-64 code from a verified domain-specific language, so it does not rely on binary lifting.
5. Conclusions and Next Steps
Macaw has been a keystone technology for us in developing binary analysis tools over a decade. Macaw’s core IR builds in multiple features that let us build rapidly and with confidence, while its symbolic execution capabilities have enabled a diverse range of testing and verification tools.
Our current focus is on improving Macaw’s code discovery capabilities. A common pain point is discovering the targets of calls through jump tables—by default, Macaw uses simple pattern-based heuristics. We have developed Macaw-Refinement (Section 2.6.1) as a more accurate alternative. However, its reliance on SMT solving can increase discovery time in a way that is unacceptable for many applications. We are currently investigating how to limit SMT calls to only the places that the pattern-based heuristic fails. Our eventual goal is to make Macaw-Refinement fast enough to be the default code discovery algorithm which would bring many more target binaries within reach.
Acknowledgements.
This work relates to Department of Navy award #N00014-22-1-2588 issued by the Office of Naval Research. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Office of Naval Research.References
- (1)
- Aho et al. (1986) Alfred V. Aho, Ravi Sethi, and Jeffrey D. Ullman. 1986. Compilers: principles, techniques, and tools. Addison-Wesley Longman Publishing Co., Inc., USA.
- Alvarez (2023) Sergi Alvarez. 2023. Libre and Portable Reverse Engineering Framework. https://www.radare.org/n/
- Anand et al. (2013) Kapil Anand, Matthew Smithson, Khaled Elwazeer, Aparna Kotha, Jim Gruen, Nathan Giles, and Rajeev Barua. 2013. A Compiler-Level Intermediate Representation Based Binary Analysis and Rewriting System. In Proceedings of the 8th ACM European Conference on Computer Systems (Prague, Czech Republic) (EuroSys ’13). Association for Computing Machinery, New York, NY, USA, 295–308. https://doi.org/10.1145/2465351.2465380
- Balakrishnan and Reps (2004) Gogul Balakrishnan and Thomas Reps. 2004. Analyzing Memory Accesses in x86 Executables. In Compiler Construction, Evelyn Duesterwald (Ed.). Springer Berlin Heidelberg, Berlin, Heidelberg, 5–23.
- Barthe et al. (2011) Gilles Barthe, Juan Manuel Crespo, and César Kunz. 2011. Relational Verification Using Product Programs. In FM 2011: Formal Methods, Michael Butler and Wolfram Schulte (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 200–214. https://doi.org/10.1007/978-3-642-21437-0_17
- Barthe et al. (2013) Gilles Barthe, Juan Manuel Crespo, and César Kunz. 2013. Beyond 2-Safety: Asymmetric Product Programs for Relational Program Verification. In Logical Foundations of Computer Science, Sergei Artemov and Anil Nerode (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 29–43. https://doi.org/10.1007/978-3-642-35722-0_3
- Bond et al. (2017) Barry Bond, Chris Hawblitzel, Manos Kapritsos, K. Rustan M. Leino, Jacob R. Lorch, Bryan Parno, Ashay Rane, Srinath Setty, and Laure Thompson. 2017. Vale: Verifying High-Performance Cryptographic Assembly Code. In Proceedings of the 26th USENIX Conference on Security Symposium (Vancouver, BC, Canada) (SEC’17). USENIX Association, USA, 917–934.
- Bosamiya et al. (2020) Jay Bosamiya, Sydney Gibson, Yao Li, Bryan Parno, and Chris Hawblitzel. 2020. Verified Transformations and Hoare Logic: Beautiful Proofs for Ugly Assembly Language. In Software Verification: 12th International Conference, VSTTE 2020, and 13th International Workshop, NSV 2020, Los Angeles, CA, USA, July 20–21, 2020, Revised Selected Papers (Los Angeles, CA, USA). Springer-Verlag, Berlin, Heidelberg, 106–123. https://doi.org/10.1007/978-3-030-63618-0_7
- Boston et al. (2021) Brett Boston, Samuel Breese, Joey Dodds, Mike Dodds, Brian Huffman, Adam Petcher, and Andrei Stefanescu. 2021. Verified cryptographic code for everybody. In CAV 2021. https://doi.org/10.1007/978-3-030-81685-8_31
- Brumley et al. (2011) David Brumley, Ivan Jager, Thanassis Avgerinos, and Edward J. Schwartz. 2011. BAP: A Binary Analysis Platform. In Computer Aided Verification, Ganesh Gopalakrishnan and Shaz Qadeer (Eds.). Springer Berlin Heidelberg, Berlin, Heidelberg, 463–469. https://doi.org/10.1007/978-3-642-22110-1_37
- Cantrill et al. (2004) Bryan M. Cantrill, Michael W. Shapiro, and Adam H. Leventhal. 2004. Dynamic Instrumentation of Production Systems. In Proceedings of the Annual Conference on USENIX Annual Technical Conference (Boston, MA) (ATEC ’04). USENIX Association, USA, 2. https://doi.org/10.5555/1247415.1247417
- Cha et al. (2012) Sang Kil Cha, Thanassis Avgerinos, Alexandre Rebert, and David Brumley. 2012. Unleashing Mayhem on Binary Code. In 2012 IEEE Symposium on Security and Privacy. 380–394. https://doi.org/10.1109/SP.2012.31
- Chipounov and Candea (2011) Vitaly Chipounov and George Candea. 2011. Enabling Sophisticated Analyses of ×86 Binaries with RevGen. In Proceedings of the 2011 IEEE/IFIP 41st International Conference on Dependable Systems and Networks Workshops (DSNW ’11). IEEE Computer Society, USA, 211–216. https://doi.org/10.1109/DSNW.2011.5958815
- Christiansen et al. (2019) David Thrane Christiansen, Iavor S. Diatchki, Robert Dockins, Joe Hendrix, and Tristan Ravitch. 2019. Dependently Typed Haskell in Industry (Experience Report). Proc. ACM Program. Lang. 3, ICFP, Article 100 (July 2019), 16 pages. https://doi.org/10.1145/3341704
- Chudnov et al. (2018) Andrey Chudnov, Nathan Collins, Byron Cook, Joey Dodds, Brian Huffman, Colm MacCárthaigh, Stephen Magill, Eric Mertens, Eric Mullen, Serdar Tasiran, Aaron Tomb, and Eddy Westbrook. 2018. Continuous Formal Verification of Amazon s2n. In Computer Aided Verification, Hana Chockler and Georg Weissenbacher (Eds.). Springer International Publishing, Cham, 430–446. https://doi.org/10.1007/978-3-319-96142-2_26
- Cohen (1993) Frederick B. Cohen. 1993. Operating System Protection through Program Evolution. Comput. Secur. 12, 6 (oct 1993), 565–584. https://doi.org/10.1016/0167-4048(93)90054-9
- David et al. (2016) Yaniv David, Nimrod Partush, and Eran Yahav. 2016. Statistical Similarity of Binaries. SIGPLAN Not. 51, 6 (jun 2016), 266–280. https://doi.org/10.1145/2980983.2908126
- David et al. (2017) Yaniv David, Nimrod Partush, and Eran Yahav. 2017. Similarity of Binaries through Re-Optimization. In Proceedings of the 38th ACM SIGPLAN Conference on Programming Language Design and Implementation (Barcelona, Spain) (PLDI 2017). Association for Computing Machinery, New York, NY, USA, 79–94. https://doi.org/10.1145/3062341.3062387
- Di Federico et al. (2017) Alessandro Di Federico, Mathias Payer, and Giovanni Agosta. 2017. rev.ng: a unified binary analysis framework to recover CFGs and function boundaries. In Proceedings of the 26th International Conference on Compiler Construction (Austin, TX, USA) (CC 2017). Association for Computing Machinery, New York, NY, USA, 131–141. https://doi.org/10.1145/3033019.3033028
- Dinaburg and Ruef (2014) Artem Dinaburg and Andrew Ruef. 2014. McSema: Static Translation of X86 Instructions to LLVM. In ReCon 2014 Conference, Montreal, Canada. https://recon.cx/2014/slides/McSema.pdf
- Dockins et al. (2016) Robert Dockins, Adam Foltzer, Joe Hendrix, Brian Huffman, Dylan McNamee, and Aaron Tomb. 2016. Constructing Semantic Models of Programs with the Software Analysis Workbench. In Verified Software. Theories, Tools, and Experiments, Sandrine Blazy and Marsha Chechik (Eds.). Springer International Publishing, Cham, 56–72. https://doi.org/10.1007/978-3-319-48869-1_5
- Dodds (2020) Joey Dodds. 2020. Announcing the ‘blst’ BLS Verification Project. https://galois.com/blog/2020/09/announcing-the-blst-bls-verification-project/
- Dullien (2020) Thomas Dullien. 2020. Weird Machines, Exploitability, and Provable Unexploitability. IEEE Transactions on Emerging Topics in Computing 8, 2 (2020), 391–403. https://doi.org/10.1109/TETC.2017.2785299
- Eagle (2011) Chris Eagle. 2011. The IDA Pro Book: The Unofficial Guide to the World’s Most Popular Disassembler. No Starch Press, USA.
- Eisenberg and Weirich (2012) Richard A. Eisenberg and Stephanie Weirich. 2012. Dependently Typed Programming with Singletons. SIGPLAN Not. 47, 12 (sep 2012), 117–130. https://doi.org/10.1145/2430532.2364522
- Fromherz et al. (2019) Aymeric Fromherz, Nick Giannarakis, Chris Hawblitzel, Bryan Parno, Aseem Rastogi, and Nikhil Swamy. 2019. A Verified, Efficient Embedding of a Verifiable Assembly Language. Proc. ACM Program. Lang. 3, POPL, Article 63 (jan 2019), 30 pages. https://doi.org/10.1145/3290376
- Gregg and Mauro (2011) Brendan Gregg and Jim Mauro. 2011. DTrace: Dynamic Tracing in Oracle Solaris, Mac OS X and FreeBSD (1st ed.). Prentice Hall Press, USA. https://doi.org/10.5555/1971960
- Hendrix et al. (2019) Joe Hendrix, Guannan Wei, and Simon Winwood. 2019. Towards verified binary raising. In Workshop on Instruction Set Architecture Specification (co-located with ITP 2019), Vol. 6.
- King (1976) James C. King. 1976. Symbolic execution and program testing. Commun. ACM 19, 7 (jul 1976), 385–394. https://doi.org/10.1145/360248.360252
- Launchbury and Peyton Jones (1995) John Launchbury and Simon L. Peyton Jones. 1995. State in Haskell. Lisp Symb. Comput. 8, 4 (dec 1995), 293–341. https://doi.org/10.1007/BF01018827
- Lee et al. (2011) JongHyup Lee, Thanassis Avgerinos, and David Brumley. 2011. TIE: Principled Reverse Engineering of Types in Binary Programs. (2 2011). https://doi.org/10.1184/R1/6469466.v1
- Naus et al. (2023) Nico Naus, Freek Verbeek, Dale Walker, and Binoy Ravindran. 2023. A Formal Semantics For P-Code. In Verified Software. Theories, Tools and Experiments.: 14th International Conference, VSTTE 2022, Trento, Italy, October 17–18, 2022, Revised Selected Papers (Trento, Italy). Springer-Verlag, Berlin, Heidelberg, 111–128. https://doi.org/10.1007/978-3-031-25803-9_7
- Protzenko et al. (2020) Jonathan Protzenko, Bryan Parno, Aymeric Fromherz, Chris Hawblitzel, Marina Polubelova, Karthikeyan Bhargavan, Benjamin Beurdouche, Joonwon Choi, Antoine Delignat-Lavaud, Cédric Fournet, Natalia Kulatova, Tahina Ramananandro, Aseem Rastogi, Nikhil Swamy, Christoph M. Wintersteiger, and Santiago Zanella-Beguelin. 2020. EverCrypt: A Fast, Verified, Cross-Platform Cryptographic Provider. In 2020 IEEE Symposium on Security and Privacy (SP). 983–1002. https://doi.org/10.1109/SP40000.2020.00114
- Protzenko et al. (2017) Jonathan Protzenko, Jean-Karim Zinzindohoué, Aseem Rastogi, Tahina Ramananandro, Peng Wang, Santiago Zanella-Béguelin, Antoine Delignat-Lavaud, Cătălin Hriţcu, Karthikeyan Bhargavan, Cédric Fournet, and Nikhil Swamy. 2017. Verified Low-Level Programming Embedded in F*. Proc. ACM Program. Lang. 1, ICFP, Article 17 (aug 2017), 29 pages. https://doi.org/10.1145/3110261
- Radare2 Book (2023) Radare2 Book. 2023. ESIL. https://book.rada.re/disassembling/esil.html
- Reid (2016) Alastair Reid. 2016. Trustworthy specifications of ARM® v8-A and v8-M system level architecture. In 2016 Formal Methods in Computer-Aided Design (FMCAD). 161–168. https://doi.org/10.1109/FMCAD.2016.7886675
- Rohleder (2019) Roman Rohleder. 2019. Hands-On Ghidra - A Tutorial about the Software Reverse Engineering Framework. In Proceedings of the 3rd ACM Workshop on Software Protection (London, United Kingdom) (SPRO’19). Association for Computing Machinery, New York, NY, USA, 77–78. https://doi.org/10.1145/3338503.3357725
- Rolles (2018) Rolf Rolles. 2018. Hex-Rays Microcode API vs. Obfuscating Compiler. https://hex-rays.com/blog/hex-rays-microcode-api-vs-obfuscating-compiler/
- Saudel and Salwan (2015) Florent Saudel and Jonathan Salwan. 2015. Triton: Concolic execution framework. In Symposium sur la sécurité des technologies de l’information et des communications (SSTIC).
- Schrijvers et al. (2008) Tom Schrijvers, Simon Peyton Jones, Manuel Chakravarty, and Martin Sulzmann. 2008. Type Checking with Open Type Functions. In Proceedings of the 13th ACM SIGPLAN International Conference on Functional Programming (Victoria, BC, Canada) (ICFP ’08). Association for Computing Machinery, New York, NY, USA, 51–62. https://doi.org/10.1145/1411204.1411215
- Schulte et al. (2022) Eric Schulte, Michael D. Brown, and Vlad Folts. 2022. A Broad Comparative Evaluation of X86-64 Binary Rewriters. In Proceedings of the 15th Workshop on Cyber Security Experimentation and Test (Virtual, CA, USA) (CSET ’22). Association for Computing Machinery, New York, NY, USA, 129–144. https://doi.org/10.1145/3546096.3546112
- Schwartz et al. (2010) Edward J. Schwartz, Thanassis Avgerinos, and David Brumley. 2010. All You Ever Wanted to Know about Dynamic Taint Analysis and Forward Symbolic Execution (but Might Have Been Afraid to Ask). In Proceedings of the 2010 IEEE Symposium on Security and Privacy (SP ’10). IEEE Computer Society, USA, 317–331. https://doi.org/10.1109/SP.2010.26
- Scott et al. (2021) Ryan Scott, Robert Dockins, Tristan Ravitch, and Aaron Tomb. 2021. Crux: Symbolic Execution Meets SMT-based Verification (Competition Contribution). https://doi.org/10.5281/zenodo.6147218
- Selfridge (2019) Benjamin Selfridge. 2019. GRIFT: A richly-typed, deeply-embedded RISC-V semantics written in Haskell. In Proc. SpISA 2019: Workshop Instr. Set Architect. Specification.
- Shoshitaishvili et al. (2016) Yan Shoshitaishvili, Ruoyu Wang, Christopher Salls, Nick Stephens, Mario Polino, Audrey Dutcher, John Grosen, Siji Feng, Christophe Hauser, Christopher Kruegel, and Giovanni Vigna. 2016. SoK: (State of) The Art of War: Offensive Techniques in Binary Analysis. In IEEE Symposium on Security and Privacy.
- Srinivasan and Reps (2015) Venkatesh Srinivasan and Thomas Reps. 2015. Synthesis of machine code from semantics. In Proceedings of the 36th ACM SIGPLAN Conference on Programming Language Design and Implementation. 596–607.
- Stephens et al. (2016) Nick Stephens, John Grosen, Christopher Salls, Andrew Dutcher, Ruoyu Wang, Jacopo Corbetta, Yan Shoshitaishvili, Christopher Kruegel, and Giovanni Vigna. 2016. Driller: Augmenting fuzzing through selective symbolic execution.. In NDSS, Vol. 16. 1–16.
- Vector35 (2023a) Vector35. 2023a. Binary Ninja. https://binary.ninja/
- Vector35 (2023b) Vector35. 2023b. Binary Ninja Intermediate Language: Overview. https://docs.binary.ninja/dev/bnil-overview.html
- Vegasena and Goodman (2019) Sai Vegasena and Peter Goodman. 2019. Binary symbolic execution with KLEE-Native. https://blog.trailofbits.com/2019/08/30/binary-symbolic-execution-with-klee-native/
- Yorgey et al. (2012) Brent A. Yorgey, Stephanie Weirich, Julien Cretin, Simon Peyton Jones, Dimitrios Vytiniotis, and José Pedro Magalhães. 2012. Giving Haskell a Promotion. In Proceedings of the 8th ACM SIGPLAN Workshop on Types in Language Design and Implementation (Philadelphia, Pennsylvania, USA) (TLDI ’12). Association for Computing Machinery, New York, NY, USA, 53–66. https://doi.org/10.1145/2103786.2103795
Appendix A Macaw Design (Additional Details)
data MemRepr (tp :: MacawType) where
BVMemRepr
:: (1 <= w)
=> NatRepr w
-> Endianness
-> MemRepr (BVType (8*w))
FloatMemRepr
:: FloatInfoRepr f
-> Endianness
-> MemRepr (FloatType f)
PackedVecMemRepr
:: NatRepr n
-> MemRepr tp
-> MemRepr (VecType n tp)
data FloatInfoRepr (fi :: FloatInfo) where
SingleFloatRepr ::
FloatInfoRepr SingleFloat
DoubleFloatRepr ::
FloatInfoRepr DoubleFloat
data Stmt arch ids where
AssignStmt
:: Assignment arch ids tp
-> Stmt arch ids
WriteMem
:: ArchAddrValue arch ids
-> MemRepr tp
-> Value arch ids tp
-> Stmt arch ids
ExecArchStmt
:: ArchStmt arch
(Value arch ids)
-> Stmt arch ids
ArchState
:: ArchMemAddr arch)
-> RegState
(ArchReg arch)
(Value arch ids)
-> Stmt arch ids
InstructionStart
:: ArchAddrWord arch
-> Text
-> Stmt arch ids
Comment
:: Text
-> Stmt arch ids
Appendix B Reopt Case Study (Additional Details)
;; Original x86-64 instruction
cmpb $0x30,0x3(%rax) ; in AT&T syntax
cmp BYTE PTR [rax+0x3],0x30 ; in Intel syntax
;; Translation without type inference
;; %t0 corresponds to %rax and is given type i64
%t1 = add i64 %t0, 0x3
%t2 = load i8, i8* inttoptr (i64 %t1 to i8*)
%t3 = icmp eq i8 %t2, 0x30
;; Translation with type inference
;; %t0 corresponds to %rax and is given type {{i8,i8,i8},{i8,i8}}*
%t1 = getelementptr {{i8,i8,i8},{i8,i8}},
{{i8,i8,i8},{i8,i8}}* %t0, i32 0, i32 1
%t2 = getelementptr {i8,i8}, {i8,i8}* %t1, i32 0, i32 0
%t3 = load i8, i8* %t2
%t4 = icmp eq i8 %t3, 0x30

### v3

@@
 We present Macaw, a modular framework that makes it possible to rapidly build reliable binary analysis tools across a range of use cases.
+Statically typed functional programming techniques are used pervasively throughout Macaw—these range from using functional optimization passes to encoding tricky architectural invariants at the type level to statically check correctness properties.
+The level of assurance that functional programming ideas afford us allow us to iterate rapidly on Macaw’s development while still having confidence that the underlying semantics are correct.
 Over a decade of development, we have used Macaw to support an industrial research team in building tools for machine code–related tasks.
@@
 This imposes a particular set of design constraints on the Macaw ecosystem, which we explain in this paper.
+The Macaw ecosystem was designed from the onset with statically typed functional programming techniques in mind.
+This can be seen in the design of Macaw’s intermediate representation (IR), which encodes information about the semantics of assembly operations and other architecture considerations at the type level.
+Macaw’s IR needs to be rigid enough to rule out obviously bad machine-code programs while flexible enough to bootstrap the semantics of existing tools which might have an ad hoc type system, such as the udis86 disassembler for x86 (Thampi, 2014) or the ARM XML specification (Reid, 2016).
 Macaw was built to support an industry research team developing novel binary analysis tools.
@@
 Section 5 concludes and discusses our future design ambitions for Macaw.
+The source code for Macaw is publicly available at https://github.com/GaloisInc/macaw.
 2. Design of the Macaw Ecosystem
@@
 This includes the core Macaw intermediate representation (IR), code discovery, optimization passes, and an ELF loader for ingesting binaries.
-Any architecture-specific functionality is encapsulated in a Macaw backend library.
-Macaw was originally designed with x86-64 support in mind, which gave rise to the first backend, Macaw-x86.
+Any architecture-specific functionality is encapsulated in a Macaw dialect library.
+Macaw was originally designed with x86-64 support in mind, which gave rise to the first dialect, Macaw-x86.
 Later, Macaw-AArch32, Macaw-PPC, and Macaw-RISCV were added.
-Each of these backends are built on top of other libraries that integrate Macaw with disassemblers for specific instruction set architectures (ISAs), and they also encode the semantics of each architecture’s instructions into a form that can be reasoned about.
-For instance, Macaw-RISCV is built on top of the GRIFT library, which provides a formal specification of the RISC-V ISA in Haskell (Selfridge, 2019).
-2.1. Integrating ISAs into Macaw
-While Macaw-Base provides the basic substrate for performing binary analysis, it is not very useful without a mechanism for representing ISA-specific details within Macaw.
-To this end, there are a suite of libraries supporting each architecture backend.
-Each Macaw-supported architecture comes with its own disassembler for parsing machine code into a form that Macaw can ingest, and it also comes with a library for representing the semantics of each instruction in a precise way.
-Having codified semantics enables both formal verification, which Macaw richly supports (see Section 3.2), and machine-code synthesis (Srinivasan and Reps, 2015), for which Macaw has experimental support.
-Incorporating a new ISA into Macaw can be a substantial amount of work.
-Throughout the years, we have experimented with a number of different approaches for implementing Macaw disassemblers and semantics, and we have made pragmatic choices over time.
-For some architecture backends, such as Macaw-RISCV, we implement disassemblers and semantics by hand.
-For other backends, we generate implementations automatically from ISA specifications.
-For instance, Macaw-AArch32’s disassembler is derived from the official ARM XML specification (Reid, 2016), which we use to compute parse tables for each instruction.
-In addition, the XML spec encodes the semantics of each instruction in ARM’s Architecture Specification Language (ASL).
-We generate Macaw-AArch32’s semantics by symbolically executing ASL expressions using Crucible (see Section 2.6).
-Each approach has its tradeoffs.
-Manually implementing disassemblers and semantics makes it easier to understand the code in isolation, but it requires more careful auditing for mistakes.
-Machine-generated code is generally quicker to develop and less error-prone, but at the expense of obfuscating the code.
-As a general principle, we prefer hand-written implementations for smaller ISAs (e.g., RISC-V) and machine-generated implementations for larger ISAs (e.g., AArch32).
-2.2. Typed Address Representations
-Macaw represents a machine address with the MemWord w type, which consists of a machine word of size w bits.
+Each of these dialects are built on top of other libraries that integrate Macaw with disassemblers for specific instruction set architectures (ISAs), and they also encode the semantics of each architecture’s instructions into a form that can be reasoned about.
+2.1. Macaw-Base: Core Data Types and Algorithms
+The Macaw-Base library is central to how Macaw works, as it defines the key data types and operations needed to perform many binary analysis tasks.
+All of Macaw’s operations revolve around an intermediate language (defined in Macaw-Base) where each IR program consists of a sequence of basic blocks.
+This IR contains all of the control flow and ISA-independent functionality needed to perform general analyses, but it is also designed to be extensible so that ISA-specific functionality can be added in particular Macaw dialects.
+We built Macaw-Base with statically typed functional programming techniques in mind.
+Various binary analysis-related algorithms can be expressed in relatively succinct fashion by applying ideas from functional programming (see Section 2.1.5).
+Moreover, we use Macaw is used in verification, so there needs to be assurance the tooling and machine code semantics are correct.
+To improve assurance, our implementation encodes information about register sizes, operation bitwidths, and architectural aspects statically using Haskell’s rich type system, thus statically checking many correctness properties of the implementation.
+2.1.1. Typed Addresses
+A key data type that is used in almost every Macaw algorithm is its notion of machine addresses.
+Macaw-Base represents a machine address with the MemWord w type, which consists of a machine word of size w bits:
 Note that w is a type-level number, which means that this convention is checked at compile-time, not runtime.
-This makes critical use of GHC’s DataKinds extension, which allows certain forms of values to be promoted to the type level (Yorgey et al., 2012).
+This makes critical use of GHC’s DataKinds extension, which allows natural numbers to be promoted to the type level using the Nat kind (Yorgey et al., 2012).
+For instance, 32-bit addresses have the type MemWord 32, 64-bit addresses have the type MemWord 64, and it is a type error to mix up the two types without performing an explicit conversion.
 The w type parameter is one example of a key design choice made throughout Macaw’s code: where possible, use GHC language extensions to encode invariants about machine code at the type level.
-This tradition was inspired by the Crucible symbolic execution library (Christiansen et al., 2019), with which Macaw integrates (see Section 2.6).
-For example, (Christiansen et al., 2019, Section 3) defines a CrucibleType data type that classifies the shape of Crucible expressions.
-Following suit, Macaw defines its own MacawType to classify the shape of machine code values—we give some selected definitions in Appendix A, Figure A.1.
-An example of this strong typing discipline being put to use is encoding the bit widths of AArch32 instructions.
+This tradition was inspired by the Crucible symbolic execution library (Christiansen et al., 2019), with which Macaw integrates (see Section 2.1.6).
+One example of where this strong typing discipline is put to use is in how Macaw encodes the bit widths of AArch32 instructions.
 AArch32 processors have two instruction sets: ARM mode, where each instruction is 32 bits, and Thumb mode, where each instruction is 16 bits.
@@
 This has caught many potential errors when developing Macaw-based tools, as ARM and Thumb instructions can often occur within the same binary.
-2.3. From Addresses to Architectures
-Just as Macaw’s address types are indexed by the number of bits w, many Macaw abstract syntax types are parameterized by an arch type parameter, which represents the processor architecture being used.
+Note that although a MemWord represents an address, it does not encode any information about where the address is in a binary (aside from a raw number), nor does it guarantee that the memory at that address will be valid.
+Macaw-Base defines additional data types to attend to these needs.
+At one level of abstraction higher than a MemWord is a MemAddr w, which consists of a region number (addrBase) and an offset into the region (addrOffset).
+Note that multiple MemAddrs can inhabit the same region:
+A MemAddr value represents a relocatable region which may (or may not) be mapped to actual memory.
+Macaw adopts the convention that the region number 0 always refers to absolute addresses.
+Other region numbers are used to represent concepts such as position-independent code.
+When Macaw encodes a binary, it needs to know information about each segment, i.e., each logically distinct sequence of a in memory, along with the corresponding addresses and permissions.
+The MemAddr type alone does not encode all this information, so Macaw-Base contains additional abstractions for this.
+First, Macaw-Base defines a MemSegment data type, which describes the overall contents of a single segment:
+Like a MemAddr, a MemSegment contains a region number (segmentBase) and an offset into that region (segmentOffset).
+Unlike MemAddr, a MemSegment also stores whether the segment has read, write, or execute permissions (segmentFlags), and it also maps each address in the segment to its underlying memory contents (segmentContents).
+The SegmentContents data type describes the mapping of addresses to memory:
+The MemChunk type characterizes whether a chunk of memory is a sequence of specific bytes (ByteRegion), a relocatable region whose contents are computed by a Relocation value (RelocationRegion), or a .bss section that is initialized with all zeroes (BSSRegion).
+The precise details of how different parts of binaries map to which MemChunks are beyond the scope of this paper.
+Macaw-Base uses the MemSegmentOff data type to describe an address that is guaranteed to be valid (i.e., point to valid memory).
+A MemSegmentOff value is simply a MemSegment plus an offset into the that segment:
+Finally, a binary is a collection of MemSegments, along with metadata describing what the entrypoint address is and which region numbers map to which segments.
+2.1.2. Typed Memory Shapes
+Another place where Macaw draws inspiration from Crucible is in how Macaw encodes the shapes of values that are stored in memory.
+Although all data in a binary can be thought of as a series of bytes, the shape of those bytes is very much dependent on the context in which it is used.
+For example, some machine instructions interpret bytes as integers, some instructions interpret bytes as floating-point values, and so on.
+It is all too easy to mix up different shapes, so Macaw encodes these shapes at the type level and parameterizes different operations based on what shapes they expect.
+Specifically, Macaw-Base defines a MacawType data type, which is shown in Figure 2.
+Note that the data constructors of MacawType are only ever meant to be used at the type level (again, using GHC’s DataKinds extension), which gives Macaw uses a pseudo-dependently typed flair.
+Aside from encoding invariants about memory shapes at the type level, it is also useful to be able to check what sort of MacawType one has at runtime.
+This is achieved via a TypeRepr object, which acts as a singleton type for MacawType (Eisenberg and Weirich, 2012).
+One can pattern-match on a TypeRepr value to determine which sort of MacawType is used at runtime.
+data TypeRepr (tp :: MacawType) where
+BVTypeRepr
+:: NatRepr n
+-> TypeRepr (BVType n)
+FloatTypeRepr
+:: FloatInfoRepr fi
+-> TypeRepr (FloatType fi)
+VecTypeRepr
+:: NatRepr n
+-> TypeRepr tp
+-> TypeRepr (VecType n tp)
+data FloatInfoRepr (fi :: FloatInfo) where
+SingleFloatRepr ::
+FloatInfoRepr SingleFloat
+DoubleFloatRepr ::
+FloatInfoRepr DoubleFloat
+2.1.3. From Addresses to Architectures
+Just as Macaw’s address types are indexed by the number of bits (w :: Nat) or a memory shape (tp :: MacawType), many Macaw abstract syntax types are parameterized by an arch type parameter, which represents the processor architecture being used.
 Doing so ensures that operations from separate architectures are not mixed up.
@@
 We refer to types like X86_64 and AArch32 as architecture extension points.
-One example of an arch-indexed type is ArchReg, which is displayed in more detail in Appendix A, Figure A.2.
+One example of an arch-indexed type is ArchReg, which is displayed in more detail in Figure 3.
 ArchReg uses GHC’s TypeFamilies extension (Schrijvers et al., 2008) to map an architecture extension point to a register data type.
@@
 A Macaw architecure uniquely determines the number of bits in a machine address, so Macaw defines a RegAddrWidth type family to bridge the gap between the former and the latter.
-This is such a commonly used type family that Macaw also provides a variety of type synonyms defined on top of RegAddrWidth for convenience, which are depicted in Appendix A, Figure A.2.
-2.4. The Macaw Intermediate Language
+This is such a commonly used type family that Macaw also provides a variety of type synonyms defined on top of RegAddrWidth for convenience, which are depicted in Figure 3.
+2.1.4. The Macaw Intermediate Language
+data Stmt arch ids where
+AssignStmt
+:: Assignment arch ids tp
+-> Stmt arch ids
+WriteMem
+:: ArchAddrValue arch ids
+-> MemRepr tp
+-> Value arch ids tp
+-> Stmt arch ids
+ExecArchStmt
+:: ArchStmt arch
+(Value arch ids)
+-> Stmt arch ids
+ArchState
+:: ArchMemAddr arch
+-> RegState
+(ArchReg arch)
+(Value arch ids)
+-> Stmt arch ids
+InstructionStart
+:: ArchAddrWord arch
+-> Text
+-> Stmt arch ids
+Comment
+:: Text
+-> Stmt arch ids
 Macaw uses a three address code–based intermediate language (Aho et al., 1986) that is centered around basic blocks (Block).
 Each basic block consists of a list of statements (Stmt) followed by a special terminator statement (TermStmt).
-These data types are given in Appendix A, Figure A.3.
-For example, the Block type is defined as follows:
+These data types are given in Figure 4.
 Each data type is parameterized by two type parameters: arch, which encodes the architecture, and ids, which denotes the set of identifiers used in assignment instructions.
@@
 The ids type parameter functions similarly to the s type parameter in Haskell’s ST s type (Launchbury and Peyton Jones, 1995).
-The control flow in Macaw is oriented around basic blocks, and each basic block has an address corresponding to the address of the first instruction.
-A disassembled machine instruction can correspond to multiple Stmts, depending on the complexity of the instruction.
+The control flow in Macaw is oriented around Blocks, and a disassembled machine instruction can correspond to multiple Stmts in a Block, depending on the complexity of the instruction.
 Note that the Stmts in two blocks are allowed to overlap.
 For example, this is necessary to support AArch32 binaries, where instructions can be in both ARM or Thumb mode depending on how they are reached.
+A single Stmt can perform an assignment (AssignStmt), write to memory (WriteMem), execute an ISA-specific statement (ExecArchStmt), or update register values (ArchState).
+The Stmt type is also used to encode metadata about which parts of a basic block correspond to the start of an instruction (InstructionStart) and any additional information that is useful for debugging purposes (Comment).
+The Stmt data type has many auxiliary types used in its data constructors, so we will define ones that are relevant to this paper as they become relevant.
 The core Macaw IR is compact and covers common operations present in multiple ISAs, which are used to represent the behavior of many instructions.
@@
 More complicated operations (i.e., ones that do not map directly to SMT) are put into syntactically separate parts of the Stmt language.
-For example, memory accesses are their own category, as they must be decomposed into SMT-friendly operations via symbolic execution (see Section 2.6).
+For example, memory accesses are their own category, as they must be decomposed into SMT-friendly operations via symbolic execution (see Section 2.1.6).
 There is also a separate category for architecture-specific extensions (e.g., data cache hints in PowerPC), for which each ISA has a small number of instructions.
-This separation between common operations and more complicated ones reflects a key Macaw design decision: the Macaw IR does not attempt to be generic enough to encode every possible instruction directly.
+Depending on where these instructions can occur in a basic block, these ISA-specific instructions are put into ArchStmt or ArchTermStmt.
+Separating common operations from more complicated ones reflects a key Macaw design decision: the Macaw IR does not attempt to be generic enough to encode every possible instruction directly.
 One benefit for this decision is that it keeps the Macaw IR relatively readable.
@@
 This gives Macaw users a high degree of control when mixing different components.
-2.5. Code Discovery
-One of the most important passes in Macaw-Base is code discovery, which aims to identify all of the functions in a binary that are reachable from a set of entry point addresses.
+2.1.5. Code Discovery
+One of the most important passes in Macaw-Base is code discovery.
+The core discovery pass identifies all of the functions in a binary that are reachable from a set of entry point addresses.
+This pass also distinguishes intra-procedural control flow like branches, jumps, and switch statements from inter-procedural control flow such as calls, tail calls, and returns.
+Code discovery is required for compositional code analysis passes and designed to work even when a binary lacks symbol tables or other static mechanisms for identifying layout.
 The core algorithm that powers code discovery is illustrated in Figure 5.
@@
 Targets of function calls are added to the exploration frontier as they are classified, which is used in subsequent iterations of the algorithm.
-2.6. Symbolic Execution via Crucible
+An example rewriting pass
+One optimization that Macaw’s rewriting pass enables is converting values to mux head normal form (MHNF), where if-then-else expressions (i.e., muxes) are raised as far up in the CFG as possible.
+As an example, consider this minimized CFG (pretty-printed for compactness):
+This CFG is not in MHNF, as it contains a mux expression (r2) that is nested within another expression’s definition (r3).
+To convert this CFG to MHNF, we want to rewrite it into something that looks more like this:
+Where 0x100a3e54 and 0x100a3e78 have folded away the constant addition of 0x100a3e50 to 0x4 and 0x28, respectively.
+This way, the instruction pointer (ip) value is directly a mux expression.
+MHNF CFGs are preferable from a code discovery perspective, as Macaw’s code discovery pass is more likely to notice branch statements when all of the branching is encoded as a top-level mux in the instruction pointer register.
+One nice consequence of Macaw being written in a functional language is that the MHNF pass can be implemented in a very simple, easy-to-follow, and relatively terse fashion.
+The MNHF pass (as well as the rest of the rewriter’s passes) are implemented in the Rewriter monad, which is a state monad that carries around the necessary state as well as the list of statements that have been rewritten so far:
+Most of the details of RewriteContext are not essential to understanding the MNHF pass, so we elide them here.
+A useful helper function that we need in order to simplify constant expressions (such as the Add expression in the example above) is rewriteApp:
+The full definition of rewriteApp is omitted for brevity, but the code above shows the essence of the function.
+Expressions that can be simplified with simple syntactic rewrites are simplified, and everything else is delegated to rewriteMhnf, the function which performs the MHNF pass:
+This function recognizes if-expressions (i.e., Muxes) that appear nested underneath bitvector additions (i.e., BVAdds) If it encounters such an if-expression, it pushes the addition down through the branches of the if-expression, simplifying constant expressions with rewriteApp along the way.
+If it encounters any other type of expression, then it assigns the expression to a fresh identifier, records the assignment in the RewriteState, and returns the assigned expression as a Value via evalRewrittenRhs.
+The MHNF rewriting pass involves a number of distinct steps: constant simplification, pattern recognition, the MNHF rewrites themselves, and appending new assignments to the current state.
+Thanks to the use of the Rewriter monad, we are able to separate the concerns of appending new assignments from the “business logic” of performing rewrites.
+Thanks to Haskell’s pattern matching, recognizing the shapes of if-expressions is straightforward.
+And thanks to classic functional programming techniques, we are able to compartmentalize the tedious (but mechanical) details of simplifying constant expressions in a single rewriteApp function.
+2.1.6. Symbolic Execution via Crucible
 A key component of the Macaw ecosystem is Macaw-Symbolic, which defines functionality for simulating Macaw IR programs with Crucible, a symbolic execution library targeting imperative code (Christiansen et al., 2019).
@@
 During simulation, Crucible will generate verification conditions that are discharged to an SMT solver.
-Besides machine code, Crucible also includes backends targeting other imperative programming languages, including C, Java, and Rust.
-Crucible’s LLVM backend (which enables reasoning about C code) has been battle-tested on a large number of verification problems from the SV-COMP verification competition (Scott et al., 2021).
+Besides machine code, Crucible also includes dialects targeting other imperative programming languages, including C, Java, and Rust.
+Crucible’s LLVM dialect (which enables reasoning about C code) has been battle-tested on a large number of verification problems from the SV-COMP verification competition (Scott et al., 2021).
 Moreover, Crucible can reason about code that mixes languages, such as C programs that use inline assembly (see Section 3.2).
-2.6.1. Symbolic Code Discovery
+Symbolic Code Discovery
 The Macaw ecosystem also includes an additional, optional method of code discovery provided by the Macaw-Refinement library.
@@
 Using Macaw-Refinement, on the other hand, allows an SMT solver to indicate the constrained set of possible jump targets to add those to the discovery frontier.
+2.2. Macaw-x86: The First Macaw Dialect
+Macaw-x86, which targets the x86-64 ISA, was the first ISA dialect that we developed alongside the Reopt tool (see Section 3.1).
+The architecture of Macaw-x86 is pictured in Figure 6.
+While this figure is specific to x86-64, it establishes a general pattern that we continued to use when developing future dialects, albeit with some notable differences (which we will address later).
+Because Macaw-x86 was our first attempt at a dialect, we did not have an existing methodology for integrating an ISA into Macaw, nor did we have a clear picture of what design choices one could make in the process of developing a dialect.
+One notable decision that we made (perhaps unconsciously) during development was to hand-write most of the code in Macaw-x86.
+This is because the Reopt tool needed to be able to run on x86-64 binaries with a relatively quick turnaround time, and as such, we could not wait until we had 100% coverage of the x86-64 ISA before demonstrating results.
+As such, we implemented support for individual x86-64 instructions on an as-needed basis, manually implementing whatever scaffolding was necessary for Macaw to reason about each instruction.
+2.2.1. Disassembling x86-64 Binaries Using Flexdis86
+The first step that Macaw-x86 performs when analyzing a binary is to diassemble it—that is, to ingest the binary’s raw bytes into a more structured representation.
+A binary’s contents can be thought of as a stream of instructions, and the role of a disassembler is to parse one instruction from the stream at a time.
+We wrote our own x86-64 disassembler library111Flexdis86 contains both an x86-64 disassembler and assembler, but we do not use the assembler from Macaw., named Flexdis86, which represents each instruction as a Haskell data type, called an InstructionInstance.
+An InstructionInstance captures all of an instruction’s relevant properties, such as its name, opcode, size, and operands (be they immediate values, registers, addresses, or otherwise).
+The main entrypoint into Flexdis86 is the disassembleInstruction function:
+The ByteReader class characterizes Haskell monads that can represent instruction streams.
+Flexdis86 is primarily meant to be used with a ByteReader instance that efficiently reads a binary from disk, but Flexdis86 also supports a simpler ByteReader instance that can be used to mock the contents of a binary for testing purposes.
+Because x86-64 is a very large ISA with well over 1,000 instructions, writing a feature-complete disassembler for x86-64 is no small feat.
+To make this task more manageable, disassembleFunction takes an NextOpcodeTable argument, which represents a lookup table that maps an instruction opcode to possible InstructionInstance values:
+To understand how NextOpcodeTable works, consider the pop instruction.
+Some versions of the pop instruction have opcodes that begin with the byte 0f, so a NextOpcodeTable would have an entry at index 0f containing a OpcodeTable that handles the remaining opcode bytes.
+This OpcodeTable would then contain another, smaller NextOpcodeTable with entries at indices a1 and a9, each containing OpcodeTableEntry values for pop.
+These table entries represent two variants of the instruction with the opcodes 0f a1 and 0f a9, respectively.
+The smaller NextOpcodeTable would also map any opcode bytes for other instructions whose opcodes begin with 0f (e.g., the addps instruction, which can use the opcode 0f 58).
+Constructing a NextOpcodeTable is by far the most tedious part of Flexdis86’s disassembler.
+To make this process less tedious, we partially automate it by leveraging the udis86 project (Thampi, 2014), which implements an x86 and x86-64 disassembler for C.
+Flexdis86 uses udis86’s XML table describing all x86-64 instructions to automatically generate a NextOpcodeTable value.
+This automation does save us quite a bit of time, but it is not perfect.
+Flexdis86’s version of the XML table makes some manual tweaks, such as adding additional instructions not covered by udis86.
+We also do not automate any of the logic for turning InstructionCandidates into full-fledged InstructionInstances, which can be surprisingly complicated due to the sheer number of instruction prefixes that x86-64 uses to distinguish different instructions.
+2.2.2. Encoding x86-64 Semantics Using Macaw-x86
+After disassembling a binary’s instructions, the next step is to reify them into a semantic representation that Macaw can reason about.
+The Macaw-x86 dialect accomplishes this by defining an InstructionDef for each x86-64 instruction, which converts an InstructionInstance from Flexdis86 into a list of Macaw basic blocks:
+We also define a small domain-specific language for defining InstructionDef values that makes it more obvious what an instruction’s semantics mean at a glance.
+As a small example, here is the complete semantics for the cbw (convert byte to word) instruction in Macaw-x86:
+These semantics use defNullary to signify that the cbw instruction does not take any operands.
+It first reads a byte from the al register (get al), uses sign extension (sext) to convert the byte into a two-byte value (n16), and finally copies the two-byte value into the ax register (ax .= ...).
+The InstructionDef DSL also includes shorthand for other common machine code idioms (immediate values, memory locations, offsets, etc.) that are not shown above.
+Unlike Flexdis86, where we made an effort to automate at least some parts of the implementation, all of the semantics implemented in Macaw-x86 are completely hand-written.
+This was largely a practical consideration, as we wanted to demonstrate results for the subset of x86-64 instructions that we encountered the most on the Reopt project, and hand-writing the semantics for these hand-picked instructions proved a fast way to accomplish this.
+Our methodology was to carefully read the Intel x86-64 ISA manual (Intel, 2024) and extract formal semantics for each instruction’s entry in the manual.
+Although the list of instructions covered by Macaw-x86’s semantics has grown over the years, it is still far from complete.
+The choice to hand-write Macaw-x86’s semantics, while expedient in the short-term, poses a challenge to long-term maintenance.
+If we use Macaw-x86 on a binary containing an instruction that is not covered by the list of covered instructions, then we must add the instruction’s semantics manually, which is a time-consuming process.
+Not only that, but this is an error-prone process, as human programmers are liable to make mistakes when translating the ISA manual’s prose.
+2.3. Macaw-PPC: A More Automated Dialect
+As we gained experience developing the Macaw-x86 dialect, it became increasingly clear that manually curating dialects for large ISAs did not scale well as well as we’d like.
+We decided that if we were to add another Macaw dialect for a big ISA, we would make a concerted effort to automate as much of the process as was reasonable.
+That opportunity came when we developed the next dialect, Macaw-PPC.
+The architecture of Macaw-PPC is pictured in Figure 7.
+Our focus with Macaw-PPC was to support both 32- and 64-bit PowerPC binaries, although we tend to exercise the 32-bit PowerPC code paths more often due to its prevalence in embedded software.
+2.3.1. Disassembling PowerPC Binaries Using Dismantle
+The first step in developing Macaw-PPC was to generate a disassembler.
+We used the udis86 disassembler in order to (partially) generate Macaw-x86’s disassembler, but because udis86 is specific is x86, we could not do something similar PowerPC.
+As a result, we developed a more general solution: a cross-architecture assembler and diassembler generation library called Dismantle.
+Dismantle would become the basis for Macaw-PPC’s disassembler, as well as the the disassembler for most Macaw dialects going forward.
+Dismantle works by consulting the data provided by LLVM’s TableGen tool (LLVM Project, 2024).
+LLVM uses TableGen as part of its code generator to describe the encodings, types, and operands of machine instructions.
+TableGen data is also ideal for generating disassemblers, so Dismantle is able to use TableGen with little or no changes required.222Typically, the only time where we need to modify the TableGen data for Dismantle’s purposes is to work around LLVM bugs.
+TableGen data files are large, and we don’t want to have to load a data file every time we disassemble a PowerPC binary, as this would be prohibitively expensive.
+Our solution is to instead load the data file only once when compiling the package.
+We use Template Haskell (Sheard and Jones, 2002) to load the PowerPC data file at compile time, parse it, and then use Template Haskell to generate code for an efficient disassembler based on the TableGen information.
+Note that we trade runtime efficiency in exchange for much more work at compile-time: it can take several minutes to compile all of Dismantle.
+This is a theme that would continue to reappear as we moved in the direction of increased automation.
+2.3.2. Semi-Automated PowerPC Semantics Using SemMC
+Just like with disassembling PowerPC, we sought to futher automate the process of encoding the semantics of PowerPC instructions into Macaw.
+Luckily for us, Heule et al.
+(2016) had recently published work on how to automatically generate formal semantics for x86-64using a techique called stratified synthesis.
+Although this work came out well after we hand-wrote our own formal semantics for x86-64, we decided to apply stratified synthesis when developing PowerPC semantics.
+We developed a library for performing this form of synthesis (for PowerPC and for other ISAs) and named it SemMC (Semantics of Machine Code).
+The SemMC library is designed to learn machine code semantics by starting with a core base set of instructions for which the semantics are hand-written.
+Using this base set, we execute SemMC on a series of fuzzer-generated machine states and instructions and use the results to learn the semantics of instructions that are not in the base set.
+To further increase our confidence that the learned semantics are correct, we run the randomly generated machine states on actual PowerPC hardware and use that as an oracle for what behavior should be considered correct.
+We repeat this process until we have a complete formal description of the ISA’s semantics.
+In this sense, SemMC’s stratified synthesis is semi-automated: some manual effort is required to describe the semantics of the base set, but everything else is fully automated.
+Much like with the semantics implementation in Macaw-x86, the semantics of SemMC base set instructions are specified using a DSL.
+For example, here are the semantics of PowerPC’s nand instruction in SemMC:
+Here, definePPCOpcode takes the instruction name and what instruction form to use.333Instruction forms describe how instructions are encoded in PowerPC.
+It is not essential to understand how instruction forms work to read this section.
+We use xform3c, which indicates an X-form instruction with three operands.
+These operands are then bound as rA, rB, and rC, which can then be used in the instruction’s definition.
+Finally, we use defLoc to store the result of performing a bitwise-AND of the values from the rS and rB operands into the location that the operand rA uses.
+SemMC converts the DSL into an intermediate language called What4.
+What4 programs are stored on disk using an S-expression–based representation so that upstream programs (e.g., Macaw-PPC) can parse and interpret the semantics they contain.
+For example, the What4 representation of the nand semantics above is the following S-expression:
+The What4 representation of SemMC semantics explicitly indicate which operands are inputs or outputs (via the in section), and it precisely captures how each output is modified when invoking the instruction.
+In addition to updating the rA operand, the What4 version also updates the instruction pointer IP, which is a lower-level detail that was elided in the DSL encoding.
+The DSL is a convenient way for humans to write What4 semantics in the base set, but SemMC also needs to be able to synthesize What4 semantics based on fuzzing.
+This is a more complicated process, as SemMC must be able to reason about how a machine’s state updates when invoking an instruction, and it must be able to do this where not all parts of the machine state are known ahead of time—that is, when some parts of the machine state are symbolic.
+To make this process simpler, SemMC leverages the Crucible (Christiansen et al., 2019) symbolic execution library.
+Crucible is primarily designed for imperative programming languages, such as C or Rust, but Crucible also works well for machine code as well.
+SemMC leverages Crucible when synthesizing What4 semantics for instructions not in the base set.
+Finally, once we have What4 semantics for all PowerPC instructions, we translate the What4 programs into Macaw CFGs using the Macaw-SemMC library.
+As was the case with Dismantle, we do not want to perform all of this conversion at runtime, as we would need to read many, many What4 S-expression files in order to look up the semantics for all of the instructions used in a typical PowerPC program.
+As such, Macaw-SemMC uses Template Haskell to look up all What4 S-expression files at compile time and generate Macaw CFG–based semantics for each PowerPC instruction, thereby avoiding the need to consult any What4 files at runtime.
+The payoff of introducing SemMC as part of the semantics translation is that we gain a significant amount of automation in the overall process, thereby avoiding some of the maintenance headaches that pervade Macaw-x86’s fully manual approach to semantics.
+There is a cost to doing so, however: SemMC’s extra layers of abstraction make it so that it approximately two minutes to compile the semantics for the entire PowerPC ISA.
+2.4. Macaw-AArch32: An Even More Automated Dialect
+Using SemMC to automate the development of Macaw-PPC offered significant advantages in how confident we were in the resulting PowerPC semantics’ correctness, and it made the code more maintainable as well.
+The SemMC approach is not foolproof, however.
+SemMC still requires human beings to curate semantics for base set instructions by hand, and while the size of the base set is smaller than the size of the overall ISA, curating the base set can still be a tedious (and potentially error-prone) process.
+When developing Macaw-AArch32, a Macaw dialect for the 32-bit ARM ISA, we had an opportunity to develop the level of automation even further.
+ARM maintains machine-readable, executable specifications of every instruction (Reid, 2016), which gives Macaw-AArch32 an official source of truth for how to disassemble and interpret any given instruction.
+Combining the offical ARM specifications with Template Haskell allows us to automate almost every aspect of Macaw-AArch32’s development.
+The architecture of Macaw-AArch32 is pictured in Figure 8.
+Like Macaw-PPC, Macaw-AArch32also relies on the Dismantle package in order to disassemble binaries.
+One notable difference in how Dismantle’s PowerPC and AArch32 dialects work is that Dismantle uses LLVM TableGen data for PowerPC, whereas it uses the ARM XML specifications for AArch32.
+Note that the constructing parse tables via the XML specifications is quite slow: it takes over an hour to build tables for the entire ISA.
+As such, we pre-compute the parse tables and store them on disk so that they can simply be loaded directly during subsequent rebuilds of Dismantle.
+There is also a SemMC dialect for AArch32, but unlike the SemMC dialect for PowerPC, it does not generate semantics for ARM instructions using stratified synthesis.
+Instead, it uses a dedicated library (named ASL-Translator) to convert the ARM XML specification’s semantics directly into the What4 intermediate language that SemMC uses.
+The XML specification is written in ASL (Reid, 2016), a custom-purpose imperative programming language with a strong types.
+As it turns out, the typing discipline that ASL uses is a natural fit for Crucible, so ASL-Translatorheavily relies on Crucible to translate the imperative, ASL-based semantics into a purely functional, What4-based encoding.
+Note that ASL-Translator is almost, but not completely, automated.
+ASL-Translator manually defines a handful of primitives to give semantics to low-level ASL operations (e.g., bitvector slicing, querying the program counter value, etc.) This is by far the trickiest part of ASL-Translator to write, and the vast majority of bug reports ultimately step from incorrectly implementing these primitives.
+Like Macaw-PPC, Macaw-AArch32’s heavy use of Template Haskell (both in ASL-Translatorand in other parts of the dialect) incurs a significant compile-time performance penalty.
+In fact, the performance penalty is even more severe for Macaw-AArch32: it takes well over 5 minutes to compile the parts of the dialect that rely on ASL-Translator-generated semantics.
+While it is likely that these times could be improved by working harder to reduce the size of the generated code, this reflects the general trend that the more one uses Template Haskell, the longer one has to wait.
+2.5. Macaw-RISCV: Revisiting a Manually-Written Dialect
+The most recently introduced Macaw dialect is Macaw-RISCV, which targets the RISC-V architecture.
+The architecture of Macaw-RISCV is pictured in Figure 9.
+In stark contrast to previous Macaw dialects, which have moved increasingly in the direction of automatically generating disassemblers and semantics for their respective ISAs, Macaw-RISCV implements its RISC-V disassembler and semantics completely by hand.
+This choice was motivated by the fact that RISC-V adheres to the reduced instruction set computer (RISC) philosophy.
+The base set of RISC-V instructions (not including extensions) only consists of 48 instructions, which are composed of a relatively small number of different opcodes.
+As such, it is viable to hand-write a Macaw dialect for RISC-V.
+2.6. The Manual-Automated Spectrum
+Each Macaw dialect exists on a spectrum of how much automation is used to implement the code supporting the dialect’s ISA, as depicted in Figure 10.
+One far end of the spectrum is hand-writing all of the code (e.g., Macaw-RISCV).
+The opposite end of the spectrum is fully automating the generation of all ISA-related code (no dialect meets this criterion exactly, but Macaw-AArch32 comes the closest).
+Having developed four Macaw dialects, we have honed some intuition regarding whether to use a hand-written approach, an automated approach, or somewhere in between when creating a new dialect.
+There are various questions one must ask to determine which approach is right for a given ISA:
+How many different instructions are there?
+It is possible to exhaustively go through all of the instructions in a small, RISC architecture, but for non-RISC architectures, the pain of going through every instruction will be more pronounced.
+How clear is the ISA specification regarding instruction semantics?
+Different ISA semantics are written with different levels of formality.
+For instance, the ARM XML specification is written with extreme formality, while the Intel x86-64 ISA manual is less so, sometimes requiring humans to interpret dense English-language prose in order to determine how certain instructions behave.
+For the latter case, a SemMC-like approach can help.
+How tolerant are you of long compile times?
+While automatically generating code with Template Haskell can increase ISA coverage and confidence in the generated code, it comes with a severe downside of noticeably increasing compile times.
+It is worth asking the question of whether this compile-time slowdown is acceptable for a given project.
 3. Case Studies
@@
 The Macaw library began its life as an internal component of the Reopt tool, and much of Macaw’s design reflects considerations that were important at the time of Reopt’s development.
+The source code for Reopt is publicly available at https://github.com/GaloisInc/reopt.
 The primary challenge in designing a tool like Reopt is going from Macaw’s IR, which only has very minimal type information associated with each basic block, to the LLVM IR, which has a comparatively richer type system.
@@
 Once the basic blocks have been recovered (c.f.
-Section 2.5), Reopt performs a demand analysis, allowing it to compute the x86-64 arguments and return registers on a per-block basis, aggregating them into register “demand sets” for entire functions, which are then synthesized into appropriate argument and return values for the whole function.
+Section 2.1.5), Reopt performs a demand analysis, allowing it to compute the x86-64 arguments and return registers on a per-block basis, aggregating them into register “demand sets” for entire functions, which are then synthesized into appropriate argument and return values for the whole function.
 At this point, we only know the size of registers, but not whether they contain integers, floating-point values, or, for those registers whose size matches that of virtual addresses for the architecture, data or code pointers.
@@
 Knowing these types, Reopt can produce the more appropriate getelementptr LLVM instruction, operating over typed pointers.
-We give an instance of such an output change in Appendix B, Figure B.1.
+We give an instance of such an output change in Figure 12.
+;; Original x86-64 instruction
+cmpb $0x30,0x3(%rax) ; in AT&T syntax
+cmp BYTE PTR [rax+0x3],0x30 ; in Intel syntax
+;; Translation without type inference
+;; %t0 corresponds to %rax and is given type i64
+%t1 = add i64 %t0, 0x3
+%t2 = load i8, i8* inttoptr (i64 %t1 to i8*)
+%t3 = icmp eq i8 %t2, 0x30
+;; Translation with type inference
+;; %t0 corresponds to %rax and is given type {{i8,i8,i8},{i8,i8}}*
+%t1 = getelementptr {{i8,i8,i8},{i8,i8}},
+{{i8,i8,i8},{i8,i8}}* %t0, i32 0, i32 1
+%t2 = getelementptr {i8,i8}, {i8,i8}* %t1, i32 0, i32 0
+%t3 = load i8, i8* %t2
+%t4 = icmp eq i8 %t3, 0x30
 3.1.2. Reopt Limitations
@@
 SAW has been successfully used to verify a variety of industrial-strength cryptographic libraries written in a mixture of C and x86-64, including the s2n Transport Layer Security (TLS) stack (Chudnov et al., 2018), the AWS LibCrypto library (Boston et al., 2021), and the BLST library (Dodds, 2020).
+The source code for SAW is publicly available at https://github.com/GaloisInc/saw-script.
 SAW ingests x86-64 machine code by leveraging Macaw to lift assembly instructions into a Crucible control-flow graph.
@@
 This is the only command that is machine code–specific, as this is where Macaw is used to lift the assembly code in increment to a CFG and pass it on to SAW’s symbolic execution engine.
-Although the llvm_verify_x86 command is specific to the x86-64 ISA, it could very easily be extended to handle other architectures as well 111For example, an llvm_verify_aarch32 command is planned: anonymized for review .
+Although the llvm_verify_x86 command is specific to the x86-64 ISA, it could very easily be extended to handle other architectures as well 444For example, an llvm_verify_aarch32 command is planned: anonymized for review .
 3.3. Other Case Studies
@@
 While this does not faithfully emulate the actual implementation of getppid, this can be an acceptable compromise to make to speed up symbolic execution, especially for WMs where the behavior of getppid is irrelevant.
+The source code for Ambient-Verifier is publicly available at https://github.com/GaloisInc/ambient-verifier.
 PATE (Patches Assured up to Trace Equivalence)
@@
 PATE then symbolically executes each CFAR in the original and patched binary, and if the patched binary fails to satisfy the frame condition inferred from the original binary, then PATE produces a summary of the conditions that lead to exhibiting different behavior.
+The source code for PATE is publicly available at https://github.com/GaloisInc/pate.
 Cerridwen
@@
 The basic blocks resulting from Macaw’s code discovery are used to compute each strand, which consists of the operations that lead to the computation of a single value within a basic block.
+The source code for Cerridwen is publicly available at https://github.com/GaloisInc/cerridwen.
 Renovate
@@
 One particular use case for Renovate is achieving binary diversity (Cohen, 1993)—that is, randomizing the layout of basic blocks to make it more resilient to attackers.
+The source code for Renovate is publicly available at https://github.com/GaloisInc/renovate.
 MCTrace
@@
 For example, MCTrace-instrumented binaries have been successfully run on bare metal using an MPC5777C PowerPC microcontroller.
+The source code for MCTrace is publicly available at https://github.com/GaloisInc/mctrace.
 Surveyor
@@
 Surveyor also supports stepping through a binary’s execution and recording a trace of symbolic events, which can be replayed later.
+The source code for Surveyor is publicly available at https://github.com/GaloisInc/surveyor.
 4. Related Work
@@
 6.
+- Heule et al.
+(2016) Stefan Heule, Eric Schkufza, Rahul Sharma, and Alex Aiken.
+2016.
+Stratified synthesis: automatically learning the x86-64 instruction set.
+In Proceedings of the 37th ACM SIGPLAN Conference on Programming Language Design and Implementation (Santa Barbara, CA, USA) (PLDI ’16).
+Association for Computing Machinery, New York, NY, USA, 237–250.
+https://doi.org/10.1145/2908080.2908121
+- Intel (2024) Intel. 2024. Intel 64 and IA-32 Architectures Software Developer’s Manual. https://www.intel.com/content/www/us/en/developer/articles/technical/intel-sdm.html
 - King (1976) James C. King. 1976. Symbolic execution and program testing. Commun. ACM 19, 7 (jul 1976), 385–394. https://doi.org/10.1145/360248.360252
@@
 - Lee et al. (2011) JongHyup Lee, Thanassis Avgerinos, and David Brumley. 2011. TIE: Principled Reverse Engineering of Types in Binary Programs. (2 2011). https://doi.org/10.1184/R1/6469466.v1
+- LLVM Project (2024) LLVM Project. 2024. TableGen Overview. https://llvm.org/docs/TableGen/
 - Naus et al.
@@
 https://doi.org/10.5281/zenodo.6147218
-- Selfridge (2019) Benjamin Selfridge. 2019. GRIFT: A richly-typed, deeply-embedded RISC-V semantics written in Haskell. In Proc. SpISA 2019: Workshop Instr. Set Architect. Specification.
+- Sheard and Jones (2002) Tim Sheard and Simon Peyton Jones.
+2002.
+Template meta-programming for Haskell.
+In Proceedings of the 2002 ACM SIGPLAN Workshop on Haskell (Pittsburgh, Pennsylvania) (Haskell ’02).
+Association for Computing Machinery, New York, NY, USA, 1–16.
+https://doi.org/10.1145/581690.581691
 - Shoshitaishvili et al.
@@
 In IEEE Symposium on Security and Privacy.
-- Srinivasan and Reps (2015) Venkatesh Srinivasan and Thomas Reps.
-2015.
-Synthesis of machine code from semantics.
-In Proceedings of the 36th ACM SIGPLAN Conference on Programming Language Design and Implementation.
-596–607.
 - Stephens et al.
@@
 1–16.
+- Thampi (2014) Vivek Thampi. 2014. udis86: Disassembler Library for x86 and x86-64. https://github.com/vmt/udis86
 - Vector35 (2023a) Vector35. 2023a. Binary Ninja. https://binary.ninja/
@@
 https://doi.org/10.1145/2103786.2103795
-Appendix A Macaw Design (Additional Details)
-data MemRepr (tp :: MacawType) where
-BVMemRepr
-:: (1 <= w)
-=> NatRepr w
--> Endianness
--> MemRepr (BVType (8*w))
-FloatMemRepr
-:: FloatInfoRepr f
--> Endianness
--> MemRepr (FloatType f)
-PackedVecMemRepr
-:: NatRepr n
--> MemRepr tp
--> MemRepr (VecType n tp)
-data FloatInfoRepr (fi :: FloatInfo) where
-SingleFloatRepr ::
-FloatInfoRepr SingleFloat
-DoubleFloatRepr ::
-FloatInfoRepr DoubleFloat
-data Stmt arch ids where
-AssignStmt
-:: Assignment arch ids tp
--> Stmt arch ids
-WriteMem
-:: ArchAddrValue arch ids
--> MemRepr tp
--> Value arch ids tp
--> Stmt arch ids
-ExecArchStmt
-:: ArchStmt arch
-(Value arch ids)
--> Stmt arch ids
-ArchState
-:: ArchMemAddr arch)
--> RegState
-(ArchReg arch)
-(Value arch ids)
--> Stmt arch ids
-InstructionStart
-:: ArchAddrWord arch
--> Text
--> Stmt arch ids
-Comment
-:: Text
--> Stmt arch ids
-Appendix B Reopt Case Study (Additional Details)
-;; Original x86-64 instruction
-cmpb $0x30,0x3(%rax) ; in AT&T syntax
-cmp BYTE PTR [rax+0x3],0x30 ; in Intel syntax
-;; Translation without type inference
-;; %t0 corresponds to %rax and is given type i64
-%t1 = add i64 %t0, 0x3
-%t2 = load i8, i8* inttoptr (i64 %t1 to i8*)
-%t3 = icmp eq i8 %t2, 0x30
-;; Translation with type inference
-;; %t0 corresponds to %rax and is given type {{i8,i8,i8},{i8,i8}}*
-%t1 = getelementptr {{i8,i8,i8},{i8,i8}},
-{{i8,i8,i8},{i8,i8}}* %t0, i32 0, i32 1
-%t2 = getelementptr {i8,i8}, {i8,i8}* %t1, i32 0, i32 0
-%t3 = load i8, i8* %t2
-%t4 = icmp eq i8 %t3, 0x30